### v1

# Opinion de-polarization of social networks with GNNs

Abstract
Nowadays, social media is the ground for political debate and exchange of opinions. There is a significant amount of research that suggests that social media are highly polarized. A phenomenon that is commonly observed is the echo chamber structure, where users are organized in polarized communities and form connections only with similar-minded individuals, limiting them- selves to consume specific content. In this paper we explore a way to decrease the polarization of networks with two echo chambers. Particularly, we observe that if some users adopt a moderate opinion about a topic, the polarization of the net- work decreases. Based on this observation, we propose an efficient algorithm to identify a good set of K users, such that if they adopt a moderate stance around a topic, the polarization is mini- mized. Our algorithm employs a Graph Neural Network and thus it can handle large graphs more effectively than other approaches
Machine Learning, ICML
1 Introduction
In the contemporary digital landscape, dominated by platforms like Instagram, Facebook, and TikTok, users find themselves immersed in an ever-expanding sea of content (Bakshy et al., 2015). Despite the sheer volume, the user’s experience usually expands to limited content. The main reason behind this phenomenon lies in the recommendation algorithms employed by these platforms. Those algorithms capture user preferences and interactions and are designed to recommend personalized content. Combined with homophily - the tendency of people to interact with similar-minded individuals. This fact leads to a significant decrease in the variety of posts that appear in the user’s feed.
This narrowing of exposure leads to a broader issue of polarization, where users organize themselves into isolated, ideologically homogeneous communities, commonly referred to as echo chambers. Within these echo chambers, individuals’ pre-existing beliefs are reinforced, often leading to extremism, fanaticism, and a diminished openness to alternative viewpoints. Addressing this challenge has become a central focus in social network analysis, with several studies proposing methods to reduce polarization, some through friend recommendation from outside the echo chamber (Tommasel et al., 2021) while others through recommendation of diverse content (Cinus et al., 2022). However, these approaches do not explicitly account for users’ opinions or quantify the level of polarization in the network. They primarily focus on diversifying content recommendations without a deeper consideration of the underlying opinion dynamics. Our main focus is on networks with two echo chambers, a phenomenon that has been observed in multiple contexts, such as politics (democrats - republicans), and COVID-19 vaccination.
In this work, we adopt the famous opinion evolution model of Friedkin and Johnsen (Friedkin & Johnsen, 1990). According to this model, the opinion that an individual adopts and expresses depends on the individual’s deeply held internal beliefs and the expressed opinions of her social environment. The details of the model are given in Section 3.
This opinion model introduces numerical values as opinions and thus it enables us to quantify the network’s polarization and try to decrease it. Moreover, the change of an opinion triggers a sequence of opinion updates for multiple users in the network, a process that is commonly referred as diffusion. Based on that, the following question rises: Which user’s opinion change, leads to the largest decrease of the total network’s polarization?
The problem that we tackle is connected to this question. Particularly, we want to identify a set of $K$ users such that when their opinion changes (in a proper way), the network’s polarization is minimized.
Finding exactly the opinion value that each of the $K$ users should have, in order to minimize polarization, is a different problem on its own. In this work, we assume that the $K$ users will adopt and express a moderate opinion. As we discuss in Section 3, the zero value opinion represents the moderate opinion for a topic.
In summary, given a limited budget $K$ we want to minimize the polarization of a social network. So we need to use the limited budget $K$ wisely to convince a proper set of users to adopt moderate opinions, such that the polarization is decreased as much as possible. This could be done either through education of users or through curated recommendations that would show to the users the opposite viewpoint about a topic, such as politics, hoping that they would eventually adopt a moderate stance. Alternatively, one might give incentives to users to adopt and express moderate opinions. In all cases the budget $K$ represents that we have limited resources for this purpose.
Answering those questions is challenging, especially in large networks where significant computation is required. Specifically, this is a combinatorial optimization problem which can’t be solved easily. The authors of (Matakos et al., 2017) proposed an algorithm to solve the problem which, although effective, it’s complexity doesn’t allow it to scale to large graphs.
2 Related Work
A significant body of research has focused on designing interventions aimed at depolarizing opinions in social networks. One common approach involves the use of content-based recommendations that expose users to diverse or opposing viewpoints in an effort to reduce ideological extremity. In (Vendeville et al., 2022) they model the Newsfeed and Wall of every user as a Markov Chain. They suggested adding curated recommendations to the newsfeeds in order to show users more diverse content. The optimal recommendation rate that maximises the average diversity according to their model, can be found by solving a system.
Another approach centers on friend recommendations, where users are encouraged to connect with individuals who hold different opinions. In (Cinus et al., 2022) they proposed a framework to asses the effects of friend-recommendation systems on the echo chamber and polarisation. They found that friend recommendation systems can strengthen echo chambers under certain conditions, while their effect might be negligible depending on the structure of the underlying network. In the context of friends recommendations, one very interesting work is (Tommasel et al., 2021). They use Graph Convolutional Networks (Kipf & Welling, 2017) along with NLP methods to construct representations of users and the echo chambers, which they later use to recommend relevant and diverse friendships.
The most closely related work to ours, is (Matakos et al., 2017), where the authors use the Friedkin and Johnsen opinion model to address the problem of minimizing overall polarization in a network. Specifically, they aim to identify a set of $K$ users whose opinions, when moderated, lead to a reduction in polarization. This paper is motivated and builds significantly on their approach, tackling the same core problem but with a method that is designed to scale more effectively to larger graphs.
3 Opinion Dynamics and Polarization Metric
We represent a social network as a graph $G(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. If node $u$ is connected with node $v$ then the tuple $(u,v)\in E$ and $w_{uv}$ is a weight associated with the edge $(u,v)$. Additionally, each node has two features. Namely, there are two opinions related with each node $i$: the internal opinion $s_{i}$ and the expressed (or external) opinion $z_{i}$, based on the Friedkin and Johnsen opinion formation model (Friedkin & Johnsen, 1990). This is a well studied and commonly used opinion model in the literature. According to it, every node $i$ in a social network maintains a fixed internal opinion $s_{i}$, while she updates her external opinion $z_{i}$ based on the following rule:
$z_{i}=\frac{s_{i}+\sum_{j\in N(i)}w_{ij}z_{j}}{1+\sum_{j\in N(i)}w_{ij}}$ (1)
The expressed opinion $z_{i}$ of node $i$ represents a compromise between the persistent value of $s_{i}$ and the expressed opinions of others to whom $i$ is connected (Bindel et al., 2015). Both $s_{i}$ and $z_{i}$ take values in the range [-1, +1] where +1 implies complete agreement while -1 complete disagreement with a topic (for example COVID19 vaccination). Intuitively, the internal opinion represents strongly held beliefs that come from the individuals education, background, origin and political orientations. On the other hand, the external opinion is the one that the individual adopts and eventually expresses.
It has been shown that if every person $i$ updates her opinion based on Equation 1, then the expressed opinions converge to a unique opinion vector $\mathbf{z}$, where each component $z_{i}$ is the opinion of user $i$ at the equilibrium (Matakos et al., 2017). The steady state vector $\mathbf{z}$ can be computed by running the evolution process until convergence or by solving the following linear system (Bindel et al., 2015):
$\mathbf{(L+I)z=s}\Leftrightarrow\mathbf{z=(L+I)^{-1}s}$ (2)
where we used the fact that $(L+I)$ is a positive definite matrix and thus invertible ($L$ is the Laplacian Matrix and $I$ is the identity matrix).
Previously, we briefly mentioned that $+1$ and $-1$ imply the complete agreement and disagreement to a topic. Other opinion values in the range $[-1,+1]$ are interpreted similarly, based on their sign, while their magnitude is proportional to their extremity. The zero opinion represents a neutral opinion. Thus, it is a natural choice to measure polarization using the polarization index that was introduced at (Matakos et al., 2017) and is defined as:
$\pi(\mathbf{z})=\frac{||\mathbf{z}||_{2}^{2}}{|V|}$ (3)
In the edge case where $\mathbf{z=0}$, all users express neutral opinion and thus the polarization index is minimized: $\pi(\mathbf{z})=0$. We divide by $|V|$ in order to make the index independent of the network’s size.
4 Network Depolarization
In this section we formalize the problem that we are dealing with and present our suggested solution. Given a social network and the opinions of it’s users, there are multiple approaches one can take to decrease the polarization index. One choice is to focus on the internal opinions $s_{i}$, and try through education to change them in a way that decreases the polarization index. Alternatively, one could focus on the external opinions and give incentives to users to start expressing opinions that will decrease the polarization. Both of these settings have been explored in (Matakos et al., 2017). In particular, they addressed the problem of identifying $K$ users such that when their opinions are set to zero the polarization index is minimized. Their experiments showed that moderating the external opinions was more effective than moderating the internal opinions, in terms of the polarization decrease. For this reason, we focus directly on the external opinions. Formally, the problem that we aim to solve is the following:
ModerateExpressed: Given a graph $G=(V,E)$, a vector of internal opinions $\mathbf{s}$, the resulting vector of expressed opinions $\mathbf{z}$, and an integer $K$, identify a set $T$ of $K$ nodes such that fixing the expressed opinions of the nodes in $T$ to 0, minimizes the polarization index $\pi$.
This is one of the problems that was introduced in (Matakos et al., 2017), where they constructed a greedy iterative algorithm to solve it. This algorithm is named GreedyExt. In the same paper, it is proved that ModerateExpressed is an NP-hard problem, and subsequently we don’t know if there exists a polynomial algorithm which can solve the problem optimally.
GreedyExt is an iterative algorithm which starts with an empty set $T^{0}$. At each step t the algorithm adds to the existing solution $T^{t-1}$ the node $v$ which when setting $z_{v}=0$ causes the largest decrease $\pi^{t-1}-\pi^{t}$ in the objective function.
In order to find the best node at each timestep the algorithm evaluates every node independently. For each node, we fix her external opinion to zero and compute the new steady state vector. The most straightforward way to do this is to run the evolution process until convergence using the Equation (1). Alternatively, the new steady state vector, after setting and fixing the external opinion of a node $i$, $z_{i}=0$, can be computed through the following process, as explained in (Matakos et al., 2017):
  1. 1.
    Set $s_{i}=0$
  2. 2.
    Update the Laplacian Matrix $\mathbf{L}$ with a row of zeros at the i-th entry
  3. 3.
    Use the Equation (2) with the new $\mathbf{s}$ and $\mathbf{L}$.
Both approaches are valid and were shown to be inefficient for large networks. In particular, solving the linear system (2) for one node involves the multiplication of an $n\times n$ matrix with an $n$-dimensional vector, which translates to $O(n^{2})$, where $n=|V|$. This process is repeated for every node and thus the complexity becomes $O(n^{3})$. Repeating $K$ times (because we are looking for $K$ nodes), leads to $O(Kn^{3})$. In (Matakos et al., 2017), they took advantage of linear algebra properties of the opinion model and were able to bring the complexity down to $O(Kn^{2})$, which, although a great improvement, is still inefficient for large graphs.
Algorithm 1 contains the pseudocode for the GreedyExt. Perhaps, one step that needs further explanation is the part where the algorithm isolates the node $v^{*}$ at each timestep. This is done in order to make sure that $z_{v^{*}}=0$, after the algorithm terminates. If we allow for $v^{*}$ to remain connected with other nodes, then in the next iteration of the algorithm, it’s external opinion will depart from zero because the node will actively participate in the opinion evolution process. As a result, only the best node $v^{*}$ of the $K$-th iteration will have $z_{v^{*}}=0$ after the algorithm terminates which is not what we aim for.
Algorithm 1 GreedyExt
  Input: Graph $G=(V,E)$, $\mathbf{s}$, $\mathbf{z}$, Maximum timesteps $K$
  Initialize solution set $T\leftarrow\emptyset$
  for $t=1\textit{ to }K$ do
     $\pi_{min}\leftarrow+\infty$
     for $node\textit{ in }V-T$ do
        $z_{node}\leftarrow 0$
        Compute the new steady-state vector $\mathbf{z_{new}}$
        $\pi\leftarrow\frac{||\mathbf{z_{new}}||_{2}^{2}}{|V|}$
        if $\pi<\pi_{min}$ then
           $\pi_{min}\leftarrow\pi$
           $v^{*}\leftarrow node$
        end if
        Reset the opinions to the original state (before the inner loop started executing)
     end for
      $T\leftarrow T\ \cup\ \left\{v^{*}\right\}$
      $z_{v^{*}}\leftarrow 0$
     Compute the new steady-state vector $\mathbf{z_{new}}$
     Update the graph with the new opinions $\mathbf{z_{new}}$
     Isolate node $v^{*}$
  end for
  Output: Solution set $T$
5 Algorithms
As we have already mentioned, GreedExt, although effectively decreases polarization, it doesn’t scale to large networks, due to its high computational complexity. Our approach is very similar to GreedyExt. In particular, we try to approximate the computationally heavy parts of the algorithm and achieve similar results in significantly less time. For this purpose we employee a Graph Neural Network (GNN), which is a neural network architecture specifically designed to work with graph data.
Graph Convolutional Network
Before we present our ML-based algorithmic approach to the ModerateExpressed problem, we introduce briefly the Graph Convolutional Network (GCN) architecture (Kipf & Welling, 2017).
Generally, GNN’s are used in order to create vector representations of nodes, edges or entire graphs depending on the underlying graph related task. We usually refer to those vector representations as embeddings. A common example of a node level task is node classification, while an example for an edge level task is link prediction. GNNs are also used for graph level tasks such as graph classification. Now, focusing on a node level task, the GNN creates the embeddings for one node based on the embeddings of its local neighborhood. A one layer GNN collects information from the direct 1-hop neighbors of a node $v$, and based on it, it constructs its embedding $h_{v}$. If the embedding must encode information that is more than 1-hop away then we have to use more layers.
GNNs are prevalent when it comes to graph representation learning, and various architectures have been proposed, such as GCN, GAT and GIN (Kipf & Welling, 2017), (Veličković et al., 2018), (Xu et al., 2018). The main difference between these architectures is the way that they aggregate the information from the neighborhood of node $v$, $N_{v}$. In this work we used the standard GCN architecture. According to it the embedding of a node $v$ at layer $k+1$ is computed as:
[…]
$h_{v}^{k+1}=\sigma(W_{k}\sum_{u\in N(v)}\frac{h_{u}^{k}}{|N(v)|}+B_{k}h_{v}^{k})$
A GCN computes the embedding of a node as the average of the embeddings of its neighbors, transformed by a non linear transformation.
ML Based Approach
Now we discuss how the GCN can help derive a method to solve efficiently the problem ModerateExpressed. Intuitively, what we want to do is to train a GCN to predict the decrease of a polarization in the network, after setting a node’s external opinion to zero. This could, theoretically, be approximated from a GCN, because the embedding of a node can encode information about the structure and the opinions of its neighborhood. Thus, a GCN should be able to predict a high decrease in the polarization, if the embedding of the node $v$ that we consider to set $z_{v}=0$, suggests that node $v$ is a central node in a significantly polarized part of the network. Based on this intuition, we continue with our algorithm which is separated into two main components: the offline part and the online part.
Offline
This part of our solution refers to the process of training the GCN to predict the decrease of the polarization after setting a node’s external opinion to zero. Let:
$Gain(v)=\pi(\mathbf{z_{old}})-\pi(\mathbf{z_{new}})$
where $\mathbf{z_{old}}$ and $\mathbf{z_{new}}$ are the vectors of external opinions before and after setting $z_{v}=0$ respectively.
In other words, we want to train a GCN to predict the $Gain$ for each node. This is a node regression task because it predicts a continuous target value. To be more precise, the GCN is responsible to generate the node embeddings. We then use those node embeddings as input to a regression head which outputs the predicted $Gain$. The regression head is usually a linear layer because the training is done in an end to end fashion and non-linearities are already applied in earlier stages of the pipeline (inside the GCN). In Section 6 we present the data that we used for the training of the GCN.
Online
Figure 1: Graphical illustration of the pipeline that finds the best node at a timestep
The online part of our algorithm is what we actually use to solve the ModerateExpressed problem. It builds upon and aims to improve, in terms of time, the previously discussed GreedyExt algorithm. The core principle behind both algorithms is iterative: at each timestep, we aim to select the best node to add to the solution set based on its contribution to the reduction of polarization in the network. However, our approach introduces a key innovation that dramatically enhances efficiency and scalability.
The fundamental difference lies in how the “best” node is identified. In the original GreedyExt algorithm, this process involved a brute-force search where, at each step, the polarization of the network was computed after setting the opinion of each node to zero individually. This required recalculating polarization multiple times, leading to substantial computational overhead, especially for large graphs. In contrast, our algorithm replaces this brute-force search with a trained neural network model, which predicts the polarization reduction directly for each node. Figure 1 summarizes the pipeline for finding the best node.
Breifly, the Graph is given as input to the GCN, which constructs one embedding for each node. Then the embeddings are fed to a regression head (a linear layer), which outputs one value per embedding. Finally, with argmax we locate the node which is the best at this timestep, according to our model.
In summary, by integrating machine learning into the node selection process, our algorithm aims to significantly improve upon the GreedyExt method, in terms of time required to construct a solution set, that is almost as good as the one that GreedyExt finds. It maintains the iterative structure, but replaces brute-force computation with a more scalable approach, reducing complexity, while preserving accuracy in the identification of key nodes. We refer to this algorithm as the GNN Algorithm in the experiments that follow. Algorithm 2 contains the pseudocode for the GNN Algorithm.
Algorithm 2 GNN Algorithm
  Input: Graph $G=(V,E)$, $\mathbf{s}$, $\mathbf{z}$, Maximum timesteps $K$, Pre-trained GNN model
  Initialize solution set $T\leftarrow\emptyset$
  for $t=1$ to $K$ do
[…]
     Select node $v^{*}\leftarrow\arg\max_{v}\textit{scores}(v)$
      $T\leftarrow T\ \cup\ \left\{v^{*}\right\}$
     Update $G\leftarrow$ set opinion of $v^{*}$ to 00 in $G$
     Compute the new equilibrium and isolate node $v^{*}$
  end for
  Output: Solution set $T$
6 Datasets and Training
We used both synthetic and real networks, which we present in this section.
Degree Corrected Stochastic Block Model
Earlier we discussed about the node regression task, which we commonly refer to as the gain regression task. In order to train the GCN for this task we need data with ground truth targets. As a first step we construct synthetic graphs with two echo chambers. The graphs are generated using a degree corrected stochastic block model (DCSBM), which was introduced in (Karrer & Newman, 2011). The difference of this model with other stochastic block models lies in the degree distribution. In the DCSBM model the degree distribution of the graph follows a power law. The details of the DCSBM model can be found in the paper (Karrer & Newman, 2011).
Figure 2: Example of graph generated with the DCSBM model
We construct 128 graphs with the DCSBM model, where each graph consists of 1000 nodes. The degrees are drawn from a power law distribution, which is common in social networks according to (Newman, 2010). Next we assign internal opinions to nodes based on their community membership. Nodes $v$ in community 1 are set $s_{v}=+1$ while nodes $u$ in community 2 are set $s_{u}=-1$. We don’t bother setting the external opinions, because their values at equilibrium don’t depend on their initial values. They depend on the vector $\mathbf{s}$ as Equation 2 suggests.
In order to use the generated graphs to train our model on the gain regression task, one more step is required. We have to compute the true $Gain$ for each node as the ground truth target value and then use it to train our model in a supervised setting. For this reason, the gain of each node has to be computed independently, as a preprocessing step for every graph of the dataset. Then the dataset can be used to train the model.
Figure 3: Retweet graph for the wiretaping scandal with 5k nodes
Our experiments suggested that it was beneficial to use simpler models in terms of the ModerateExpressed problem. The more complex models, although they performed better in terms of the gain regression task, they didn’t make as good node choices in the ModerateExpressed problem. This is probably because the more complex models could learn very well the properties of the graphs generated from the DCSBM model. In other words, we suspect that the complex models didn’t overfit the dataset but they overfitted the global properties of the synthetic graphs, which are naturally different than real graphs. However, at the end of the day we are interested in the ModerateExpressed problem and not in the gain regression. Consequently, larger loss in the gain regression task is acceptable, as long as the ranking of the nodes is correct. We are not interested in the exact value of the gain. We only want to be able to tell which node is the best choice at each timestep. Finally, a simpler model has an extra advantage in terms of the time required for the algorithm to complete. As a result, for our experiments we used a simple GCN with 2 layers where each layer consisted of 16 hidden units.
Real Networks
We evaluated the GNN Algorithm in both real and synthetic networks. Here we present the real networks that we used. The experimental results are discussed in the next section.
Political Books
This dataset consists of political books that were sold on Amazon. The edges between political books represent frequent co-purchasing. The books are labeled as Conservative, Liberal and Neutral. The opinion mapping that we use is the following: Conservative: 1, Neutral: 0, Liberal: -1. The reason why we choose these 3 values is because we have no means to quantify numerically for each node what their exact opinion is. One alternative is to randomly sample opinions in the negative side of the spectrum of Liberal nodes, and in the positive side of the spectrum for the Conservative nodes. Finally, for the neutral nodes we could sample opinions around zero. There is no big difference in terms of our methodology. Of course, the resulting polarisation of the network and the final selected nodes by the algorithms would be different as their selection or not from the algorithms, depends both on their opinion and graph structure. But we are examining whether the GNN Algorithm that we propose makes equivalent choices to the GreedyExt algorithm, so the starting point of the network doesn’t really make any difference. In both cases the GreedyExt algorithm will make its greedy choices, and our algorithm should present similar behavior. For our experiments we decided to stick with the +1,0, -1 mapping to align with the equivalent experiments in (Matakos et al., 2017). The polarisation is $\pi=0.107$. It consists of 105 nodes and 441 edges.
LiveJournal
LiveJournal is a free online blogging community where users declare friendship to each other. LiveJournal also allows users to form a group, to which other members can then join. Such user-defined groups are considered as ground-truth communities. This dataset was introduced in (Yang & Leskovec, 2012) and is made public from the authors (SNAP Datasets). It consist of multiple communities. Because of our interest in the dual echo chamber structure, a network with two communities was required. For this reason, we started with the 5000 best communities according to the original paper (Yang & Leskovec, 2012), where they define what makes a community “good”. Then we sorted them based on their length and constructed the graph for the two largest communities. The resulting graph is depicted in Figure 4 and clearly has the dual echo chamber structure. The opinions are assigned based on community membership.
Figure 4: LiveJournal Graph
Twitter Wiretaping Scandal
This dataset contains a snapshot of the political discussion on Twitter regarding the Greek Wiretaping Scandal, which was revealed in 2022. The detailed description of this dataset is given in (Dimitriadis et al., 2024). Briefly, it contains multiple tweets with hashtags and keywords relevant to the scandal. Using this dataset we construct the retweet graph. To infer user opinions, we adopt a straightforward heuristic. We count the number of politicians that a user follows and assume that the party she supports is the one that she follows the most politicians from. In case of a draw we assume she doesn’t support any party. Finally, we set her opinion to $s_{i}=+1$ if she supports a party that is left on the political spectrum and $s_{i}=-1$ otherwise. Because the retweet graph consists of 44k nodes, we work with a smaller sample of 5k nodes which is depicted in Figure 3. Due to its high density, this graph doesn’t present as high polarization as the other graphs.
7 Experiments
In this section we present the experimental results of our algorithm and compare it to the GreedyExt. For all experiments we set $K$ to be small, as allowing for large values makes the problem trivial. For example if we set $K$ to be equal to approximately 90% of the graph’s size, then we wouldn’t need to make sophisticated node choices, as most of the nodes would be by default set to zero and the polarization index would have been negligible. We limit $K$ to approximately 10% of the network’s size.
Additionally, in order to demonstrate that the problem is not trivial, we used a random algorithm as a baseline. The random algorithm picks $K$ nodes randomly and sets their external opinions to zero. Similar performance as the GreedyExt implies that the problem is trivial.
Figure 5: Polarization vs K for the Books Dataset
Polarization
Figure 5 shows how the polarization changes on each step of the algorithms, for the Political Books dataset. For example the green point at $k=4$ denotes the polarization of the network after the GNN Algorithm has set to zero the first 4 nodes (out of $K$). We observe the following points. First of all, the random algorithm fails to decrease the polarization as much as the other algorithms, a fact which implies that the problem is not trivial. Additionally, the GNN Algorithm follows GreedyExt closely, which means that it makes equivalent choices at each timestep.
The results of our experiments for the real networks are summarized in Figure 6. According to this figure, the GNN Algorithm always achieves similar polarization to GreedyExt. The organe bar is the initial polarisation for each dataset, while the blue and the green bars show the polarization achieved from GreedyExt and GNN Algorithm respectively, after setting $K$ nodes to zero.
Figure 6: Polarization of each algorithm per dataset
Time
So far, we have only discussed how the two algorithms compare, in terms of the final polarization that they achieve. As we have already mentioned, the GreedyExt algorithm, is not efficient for large graphs, because of its high computational complexity. Particularly, the evaluation of every node independently at each timestep increases the complexity dramatically. In contrast, the GNN Algorithm replaces this brute-force style approach with a forward pass on a relatively small neural network, followed by an argmax operation. Figure 7 shows the required time for each algorithm to complete for the datasets LiveJournal and Wiretaping. For this experiment we set $K=50$.
Figure 7: Required time per algorithm for the LiveJournal and Wiretaping datasets
As the picture shows, the GNN Algorithm outperforms the GreedyExt in terms of time. Similar results hold for the other datasets as well. Finally, Figure 8 shows the required time for each algorithm to complete on graphs with various sizes. As we can see, the required time for the GreedyExt algorithm grows very quickly with respect to the graph size. In contrast the time to complete for the GNN Algorithm grows much slower, with respect to the size of the graph.
Figure 8: Required time per algorithm for graphs with various sizes
Moreover, the time advantage of the GNN Algorithm against the GreedyExt, is larger, for larger graphs. In other words, it increases with respect to the size of the underlying graph. In the examples that we have shown where large graphs were involved, the LiveJournal and the Wiretaping datasets, the GNN Algorithm is ×$\times 10$ and ×$\times 16$ times faster respectively.
8 Conclusion
In this work, we proposed an alternative approach to address the ModerateExpressed problem, originally introduced in (Matakos et al., 2017). One of the key limitations of their solution is its lack of scalability, as their algorithm becomes increasingly slow, when applied to larger graphs. Our method leverages Graph Neural Networks (GNNs), specifically a 2-layer Graph Convolutional Network (GCN), to generate node embeddings. Instead of using brute force evaluation, we apply a simple argmax function to identify the node with the highest score. In terms of polarization, our approach achieves results almost identical to those of the original algorithm. However, the main advantage of our method is its ability to scale efficiently to larger graphs due to the GNN architecture. Our results show that the GNN Algorithm dramatically improves the required time to solve the problem and thus it is a great choice when dealing with large graphs. An interesting future direction is to explore the submodularity of the problem ModerateExpressed and quantify the suboptimality for greedy algorithms like the ones we have discussed.