### v2

# Opinion de-polarization of social networks with GNNs

Abstract
Nowadays, social media is the ground for political debate and exchange of opinions. There is a significant amount of research that suggests that social media are highly polarized. A phenomenon that is commonly observed is the echo chamber structure, where users are organized in polarized communities and form connections only with similar-minded individuals, limiting them- selves to consume specific content. In this paper, we explore a way to decrease the polarization of networks with two echo chambers. Existing work based on popular opinion diffusion models in the literature suggests that, convincing a subset of “key” users to adopt a moderate opinion about a topic, can help reduce the polarization of the entire network. Based on this result, we propose an efficient algorithm to identify a good set of K users such that if they adopt a moderate stance around a topic, the polarization is minimized. Unlike existing work that requires analytically solving a large model iteratively, we propose an algorithm based on Graph Neural Networks (GNN) that (i) can significantly reduce the decision time, (ii) achieve close-to-oracle selection accuracy, and (iii) be robust to the choice of opinion diffusion model.
Machine Learning, ICML
1 Introduction
In the contemporary digital landscape, dominated by platforms like Instagram, Facebook, and TikTok, users find themselves immersed in an ever-expanding sea of content (Bakshy et al., 2015). Despite the immense volume of information available on almost every topic, the user’s experience usually expands to limited content. This is for a number of reasons. Homophily, the tendency of people to interact with similar-minded individuals, leads users to form friendships (both online and offline) with similar-minded peers. Psychological phenomena such as biased assimilation and selective exposure can further obstruct a user’s exposure to differing opinion content (Greitemeyer et al., 2009) (Frey, 1986). Last but not least, ubiquitous recommendation algorithms that largely guide a user’s attention are designed to recommend content the user will probably agree with, leading to ever increasing polarization on many sensitive topics.
This narrowing of exposure leads to a broader issue of polarization, where users organize themselves into isolated, ideologically homogeneous communities, commonly referred to as echo chambers. Within these echo chambers, individuals’ pre-existing beliefs are reinforced, often leading to extremism, fanaticism, and a diminished openness to alternative viewpoints. Addressing this challenge has become a central focus in social network analysis, with several studies proposing methods to reduce polarization, some through friend recommendation from outside the echo chamber (Tommasel et al., 2021) while others through recommendation of diverse content (Vendeville et al., 2022). Most such works directly try to identify what content or friendship to recommend to a target user. For instance, a framework that predicts the probability of a new friendship given the users’ previous tweets and interactions is proposed in (Tommasel et al., 2021). Such probabilities are used to recommend friendships from opposing echo chambers that are likely to be formed and thus lead to a decrease in the modularity of the network.
Polarization is interwoven with the opinions of the users of a network, which are often represented numerically. In this work, we adopt one of the most widely used opinion evolution models, known as the Friedkin and Johnsen (FK) model (Friedkin & Johnsen, 1990) (Biondi et al., 2023). According to this model, the opinion that an individual adopts and expresses is represented numerically and depends on the individual’s deeply held internal beliefs and the expressed opinions of her social environment. The details of the model are given in Section 3.
Given this interdependence of the opinions of different users that the FK model suggests, the change in one’s opinion triggers a sequence of opinion updates for multiple users in the network, a process that is commonly referred to as diffusion. Based on that, the following question rises: Which user’s opinion change, leads to the largest decrease of the total network’s polarization?
The problem that we tackle is connected to this question. Specifically, the goal of this work is to accurately and efficiently identify a subset of K users such that their adoption of moderate opinion (moderation) minimizes polarization [^1] How this opinion moderation will take place is orthogonal to our work, and could involve education, incentives (e.g. monetary) and other means, as suggested by (Matakos et al., 2017). Without loss of generality, we will assume that “moderating a user” means setting the level of her expressed opinion to 0 which represents the neutral opinion around a topic, as discussed in Section 3. This is a non-trivial task, because multiple factors including the extremity of the user’s opinion and her friendships must be taken into account when examining whether the user is a good candidate to be moderated.
This problem has been considered in the seminal work of (Matakos et al., 2017), where it is proved that it is an NP-hard problem. The same work proposes an iterative greedy algorithm which selects the node whose moderation minimizes polarization at each timestep, and updates the network accordingly. The greedy algorithm is further compared with reasonable heuristics for the problem and outperforms them. The details of this algorithm are presented in Section 5, but its main drawback is its lack of scalability to large networks, as discussed by the authors.
To this end, this work proposes a novel method which utilizes modern Machine Learning methods to enable efficient identification of key nodes to moderate. We propose a similar greedy algorithm which integrates a Graph Neural Network to simplify the most computationally expensive steps, which are essential for selecting the optimal node at each timestep. Compared to the algorithm introduced in (Matakos et al., 2017), our approach constructs the solution set dramatically faster, while incurring only a minimal difference in the final polarization of the network.
The contributions of this work are:
  • We introduce a Graph Neural Network architecture which generates a score that reflects the effectiveness of a user’s moderation in reducing overall polarization, under some opinion model.
  • Our algorithm utilizes this GNN to solve efficiently the NP-hard problem of choosing a subset of K nodes whose moderation minimizes polarization. We demonstrate up to 16x speedup while maintaining 95% of the performance, compared to previous approaches.
  • We evaluate our model on both synthetic and real-world data, demonstrating its robustness in various scenarios.
2 Related Work
A significant body of research has focused on designing solutions aimed at depolarizing opinions in social networks. One common approach involves the use of content-based recommendations that expose users to diverse or opposing viewpoints in an effort to reduce ideological extremity. In (Vendeville et al., 2022) a theoretical model that describes how content flows through the platform is used. They suggested adding curated recommendations to the users’ newsfeeds in order to show users more diverse content. The optimal recommendation rate of curated content that maximises the average newsfeed’s diversity according to their model, is found via a quadratic program with linear constraints.
Another approach centers on friend recommendations, where users are encouraged to connect with individuals who hold different opinions. A framework to asses the effects of friend-recommendation systems on the echo chamber and polarisation is proposed in (Cinus et al., 2022). They found that friend recommendation systems can strengthen echo chambers under certain conditions, while their effect might be negligible depending on the structure of the underlying network. In the context of friends recommendations, one very interesting work is (Tommasel et al., 2021). They use Graph Convolutional Networks (Kipf & Welling, 2017) along with NLP methods to construct representations of users and the echo chambers, which they later use to recommend relevant and diverse friendships. While our work also utilizes Graph Convolutional Networks, it is different from (Kipf & Welling, 2017) which constructs a friendhsip recommendation system that recommends friends from outside the target user’s echo chamber, and makes no use of an opinion model.
Unlike those works, our approach attempts to tackle the specific problem setup introduced in (Matakos et al., 2017), where under the FK opinion model and a polarization metric, the authors try to select $K$ users, which when moderated lead to the smallest polarization. This paper is motivated and builds significantly on their approach, tackling the same core problem but with a method that is designed to scale more effectively to larger graphs.
3 Opinion Dynamics and Polarization Metric
We represent a social network as an undirected graph $G(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. If node $u$ is connected with node $v$ then the tuple $(u,v)\in E$ and $w_{uv}$ is a weight associated with the edge $(u,v)$. Additionally, each node has two features. Namely, there are two opinions related with each node $i$: the internal opinion $s_{i}$ and the expressed (or external) opinion $z_{i}$, based on the Friedkin and Johnsen opinion formation model (Friedkin & Johnsen, 1990). This is a well studied and commonly used opinion model in the literature (Biondi et al., 2023).
According to it, every node $i$ in a social network maintains a fixed internal opinion $s_{i}$ [^2] Fixed internal opinion does not mean that the user doesn’t change her beliefs and only is affected in terms of what she expresses. Instead, it tries to capture some biases and the various psychological factors, background, education, origin etc. that make some users more or less “pliable” to peers’ opinions, while she updates her external opinion $z_{i}$ based on the following rule:
$z_{i}=\frac{s_{i}+\sum_{j\in N(i)}w_{ij}z_{j}}{1+\sum_{j\in N(i)}w_{ij}}$ (1)
The expressed opinion $z_{i}$ of node $i$ represents a compromise between the persistent value of $s_{i}$ and the expressed opinions of others to whom $i$ is connected (Bindel et al., 2015). Both $s_{i}$ and $z_{i}$ take values in the range [-1, +1] where +1 implies complete agreement while -1 complete disagreement with a topic (for example COVID19 vaccination). The expressed opinion is the actual opinion of the user and according to the FK model it depends both on her “environment” and internal biases that she has.
It has been shown that if every person $i$ updates her opinion based on Equation 1, then the expressed opinions converge to a unique opinion vector $\mathbf{z^{ss}}$, where each component $z_{i}^{ss}$ is the opinion of user $i$ at the equilibrium (Matakos et al., 2017). The steady state vector $\mathbf{z^{ss}}$ can be computed by running the evolution process until convergence or by solving the following linear system (Bindel et al., 2015):
$\mathbf{(L+I)z^{ss}=s}\Leftrightarrow\mathbf{z^{ss}=(L+I)^{-1}s}$ (2)
In the above $L$ is the weighted Laplacian matrix defined as:
$\mathbf{L}=\begin{cases}\sum_{j\in N(i)}w_{ij},&\text{if }i=j\\ -w_{ij}&\text{if }i\neq j\text{ and }(i,j)\in E\\ 0&\text{otherwise}\end{cases}$
and $\mathbf{I}$ is the identity matrix. $\mathbf{L+I}$ is invertible as a positive definite matrix. Notice that based on Equation 2 the steady state vector $\mathbf{z^{ss}}$ depends only on the internal opinions $\mathbf{s}$ and the Laplacian matrix $\mathbf{L}$ of the underlying graph.
Previously, we briefly mentioned that $+1$ and $-1$ imply the complete agreement and disagreement to a topic. Other opinion values in the range $[-1,+1]$ are interpreted similarly, based on their sign, while their magnitude is proportional to their extremity. The zero opinion represents a neutral opinion. Without loss of generality and in order to facilitate comparison between our results and the results of (Matakos et al., 2017), we use the same polarization metric, known as polarization index. Polarization index is defined as:
$\pi(\mathbf{z^{ss}})=\frac{||\mathbf{z^{ss}}||_{2}^{2}}{|V|}$ (3)
In the edge case where $\mathbf{z^{ss}=0}$, all users express neutral opinion and thus the polarization index is minimized: $\pi(\mathbf{z^{ss}})=0$. We divide by $|V|$ in order to make the index independent of the network’s size. Nevertheless, instead of using the $L_{2}$ norm, we could have used other norms of $\mathbf{z}$, such as the $L_{1}$ norm, in order to measure polarization.
4 Network Depolarization
In this section we formalize the problem that we are dealing with and present our suggested solution. Given a social network and the opinions of it’s users, there are multiple approaches one can take to decrease the polarization index. One choice is to focus on the internal opinions $s_{i}$, and try through education to change them in a way that decreases the polarization index. Alternatively, one could focus on the external opinions and give incentives to users to start expressing opinions that will decrease the polarization. Both of these settings have been explored in (Matakos et al., 2017). In particular, they addressed the problem of identifying $K$ users such that when their opinions are set to zero the polarization index is minimized. Their experiments showed that moderating the external opinions was more effective than moderating the internal opinions, in terms of the polarization decrease. For this reason, we focus directly on the external opinions. Formally, the problem that we aim to solve is the following:
ModerateExpressed: Given a graph $G=(V,E)$, a vector of internal opinions $\mathbf{s}$, the resulting steady state vector of expressed opinions $\mathbf{z}$, and an integer $K$, identify a set $T$ of $K$ nodes such that fixing the expressed opinions of the nodes in $T$ to 0, minimizes the polarization index $\pi$.
In simple words, this problem requires us to:
  1. 1.
    Select $K$ nodes and put them in a set $T$
  2. 2.
    For each node $i\in T$ set $z_{i}=0$
  3. 3.
    Compute the new steady state opinion vector $\mathbf{z^{ss}}$, while maintaining $z_{i}=0,\forall i\in T$ (opinions of nodes in $T$ are “anchored” to zero)
  4. 4.
    The lowest $\pi(\mathbf{z^{ss}})$, the better our solution set $T$ is.
Step 3 requires computing the new steady state vector $\mathbf{z^{ss}}$. The most straightforward way to do this is to run the evolution process until convergence using the Equation (1). It is important to manually prohibit nodes in $T$ to depart from zero while running the evolution process.
Alternatively, the new steady state vector, after setting and fixing the external opinions $z_{i}=0,\forall i\in T$, can be computed through the following process, as explained in (Matakos et al., 2017):
  1. 1.
    Set $s_{i}=0,\forall i\in T$
  2. 2.
    Update the Laplacian Matrix $\mathbf{L}$ with a row of zeros at the $i$-th row, $\forall i\in T$
  3. 3.
    Use the Equation (2) with the new $\mathbf{s}$ and $\mathbf{L}$ to get the new $\mathbf{z^{ss}}$.
Next, we will provide a mathematical formulation of the problem ModerateExpressed, but in order to do so, we have to present some notation first:
Let $\mathbf{x}$ be a binary vector (it’s components are either 1 or 0), where $x_{i}=1$ iff node $i\in T$. Furthermore let $\mathbf{D_{x}}$ be the diagonal matrix with the components of $\mathbf{x}$ on it’s main diagonal. Finally $\mathbf{\overline{x}}$ is the complementary vector of $\mathbf{x}$ with components:
$\overline{x_{i}}=\begin{cases}1,\quad\text{if }x_{i}=0\\ 0,\quad\text{otherwise}\end{cases}$
Then we can compute the new $\mathbf{z^{ss}}$ which results from moderating the nodes in $T$ as:
$\mathbf{z^{ss}}=(\mathbf{D_{\overline{x}}}\mathbf{L+I})^{-1}\cdot(\mathbf{s-(s\odot x)})$ (4)
where $\odot$ is the Hadamard product (element-wise multiplication) between $\mathbf{s}$ and $\mathbf{x}$. The matrix $\mathbf{D_{\overline{x}}}$ is used to fill with zeros the rows dictated by the vector $\mathbf{x}$, while the $\mathbf{s-(s\odot x)}$ part of expression (4) is responsible to zero out the corresponding components of $\mathbf{s}$. As we can see $\mathbf{z^{ss}}$ is a function of $\mathbf{x}$ and thus, the polarization index $\pi$ (Eq. 3) is a function of $\mathbf{x}$.
Finally, let:
$f(\mathbf{x})=\frac{||(\mathbf{D_{\overline{x}}}\mathbf{L+I})^{-1}\cdot(\mathbf{s-(s\odot x)})||_{2}^{2}}{|V|}$ (5)
Here $f$ is simply the polarization index $\pi$ of the new steady state vector, expressed as a function of $\mathbf{x}$
With that in mind we are now ready to define formally the problem ModerateExpressed:
$\text{minimize}_{\mathbf{x}}\ f(\mathbf{x})$
subject $\text{subject to}\ \sum_{i=1}^{|V|}x_{i}=K$
With this formulation, we are essentially looking for the binary vector $\mathbf{x}$ with exactly $K$ non zero components, such that the quantity $f\mathbf{(x)}$ in Equation 5 is minimized.
The problem ModerateExpressed is one of the problems that was introduced in (Matakos et al., 2017), and it was proven that it is an NP-hard problem. Subsequently, we don’t know if there exists a polynomial algorithm which can compute an optimal solution. The authors constructed a greedy iterative algorithm which is named GreedyExt in order to solve the problem. Next we present this algorithm.
GreedyExt is an iterative algorithm which starts with an empty set $T^{0}$. At each step $t$ the algorithm adds to the existing solution $T^{t-1}$ the node $v$ which when setting $z_{v}=0$ causes the largest decrease $\pi(\mathbf{z^{ss}_{t-1}})-\pi(\mathbf{z^{ss}_{t}})$ in the polarization index.
In order to find the best node at each timestep the algorithm evaluates every node independently. For each node, we fix her external opinion to zero and compute the new steady state vector along with the decrease in the polarization. After evaluating every node, the one that caused the greatest decrease in the polarization is added to the solution set $T$.
Algorithm 1 contains the pseudocode for the GreedyExt. Perhaps, one step that needs further explanation is the part where the algorithm isolates the node $v^{*}$ at each timestep. This is done in order to make sure that the selected node remains “anchored” to zero throughout the entire algorithm execution. Skipping this step would allow the nodes that have been set to zero in previous timesteps to participate in the evolution process of future timesteps. As a result these nodes would depart from zero which is not what we aim for [^3] In fact this step is required in the case where the new steady state vector $\mathbf{z^{ss}}$ is computed via running the evolution process until convergence, using the Equation (1). In the case where the analytical solution from (2) is used, this step is implicitly taken care of from setting the row of $\mathbf{L}$ for the selected node to zero .
This algorithm, although efficient in finding a good solution, suffers from high computational complexity. Computing the new polarization index after the moderation of each node independently is quite expensive. Specifically, solving the linear system (2) for one node involves the multiplication of an $n\times n$ matrix with an $n$-dimensional vector, which translates to $O(n^{2})$, where $n=|V|$. This process is repeated for every node and thus the complexity becomes $O(n^{3})$. Repeating $K$ times (because we are looking for $K$ nodes), leads to $O(Kn^{3})$.
Fortunately, the authors of (Matakos et al., 2017) observed that for the specific evolution model, $\mathbf{z^{ss}}$ can be computed more efficiently, bringing the complexity of GreedyExt down to $O(Kn^{2})$. Despite a clear improvement in speed, this complexity can still be prohibitively even for moderate values of $K$ and n especially, as we’ll also show in our experiments. Furthermore, variations of the FK model (Biondi et al., 2023) (as well as other opinion evolution models) abound in related work, and there’s no guarantee that such analytical shortcuts (as the one that brought the complexity down to $O(Kn^{2})$) exist for these models.
This work aims to provide a solution to both of these issues. Regarding the complexity of GreedyExt we suggest to incorporate Machine Learning to produce accurate predictions for the decrease of the polarization after moderating a node. With such method, we would be able to replace the computationally expensive part of GreedExt, which has to compute the steady state vector after moderating every node individually. Instead of doing that, with our approach we can select a node at each timestep based on our model’s predictions, significantly speeding up the entire process. Furthermore, we claim that our ML model that produces those predictions is in some way ”opinion-evolution-model-agnostic”. By that, we mean that we could train the ML model on top of different opinion evolution models and predict the corresponding decrease in the polarization. This way we have an efficient version of GreedyExt which is no longer tied to specific opinion-evolution-model’s optimization margins.
Algorithm 1 GreedyExt
  Input: Graph $G=(V,E)$, $\mathbf{s}$, $\mathbf{z}$, Maximum timesteps $K$
  Initialize solution set $T\leftarrow\emptyset$
  for $t=1\textit{ to }K$ do
     $\pi_{min}\leftarrow+\infty$
     for $node\textit{ in }V-T$ do
        $z_{node}\leftarrow 0$
        Compute the new steady-state vector $\mathbf{z_{new}}$
        $\pi\leftarrow\frac{||\mathbf{z_{new}}||_{2}^{2}}{|V|}$
        if $\pi<\pi_{min}$ then
           $\pi_{min}\leftarrow\pi$
           $v^{*}\leftarrow node$
        end if
        Reset the opinions to the original state (before the inner loop started executing)
     end for
      $T\leftarrow T\ \cup\ \left\{v^{*}\right\}$
      $z_{v^{*}}\leftarrow 0$
     Compute the new steady-state vector $\mathbf{z_{new}}$
     Update the graph with the new opinions $\mathbf{z_{new}}$
     Isolate node $v^{*}$
  end for
  Output: Solution set $T$
5 Proposed Solution
In the end of the previous section we briefly introduced the idea of incorporating a Machine Learning model to enable efficient evaluation of candidate solution nodes. In this section we discuss the details of our suggested approach, which utilizes a Graph Neural Network (GNN). A GNN is a Neural Network architecture that is specifically designed to work with graph data.
Graph Convolutional Network
Before we present our ML-based algorithmic approach to the ModerateExpressed problem, we introduce briefly the Graph Convolutional Network (GCN) architecture (Kipf & Welling, 2017).
Generally, GNN’s are used in order to create vector representations of nodes, edges or entire graphs depending on the underlying graph related task. We usually refer to those vector representations as embeddings. A common example of a node level task is node classification, while an example for an edge level task is link prediction. GNNs are also used for graph level tasks such as graph classification. Now, focusing on a node level task, the GNN creates the embeddings for one node based on the embeddings of its local neighborhood. A one layer GNN collects information from the direct 1-hop neighbors of a node $v$, and based on it, it constructs its embedding $h_{v}$. If the embedding must encode information that is more than 1-hop away then we have to use more layers.
GNNs are prevalent when it comes to graph representation learning, and various architectures have been proposed, such as GCN, GAT and GIN (Kipf & Welling, 2017), (Veličković et al., 2018), (Xu et al., 2018). The main difference between these architectures is the way that they aggregate the information from the neighborhood of node $v$, $N_{v}$. In this work we used the standard GCN architecture. According to it the embedding of a node $v$ at layer $k+1$ is computed as:
[…]
$h_{v}^{k+1}=\sigma(W_{k}\sum_{u\in N(v)}\frac{h_{u}^{k}}{|N(v)|}+B_{k}h_{v}^{k})$
A GCN computes the embedding of a node as the average of the embeddings of its neighbors, transformed by a non linear transformation.
GCN for ModerateExpressed
Now we discuss how the GCN can help derive a method to solve efficiently the problem ModerateExpressed. Intuitively, what we want to do is to train a GCN to predict the decrease of the polarization of a network, after setting a node’s external opinion to zero. This could, theoretically, be approximated from a GCN, because the embedding of a node can encode information about the structure and the opinions of its neighborhood. Thus, a GCN should be able to predict a high decrease in the polarization, if the embedding of the node $v$ that is being moderated, suggests that node $v$ is a central node in a significantly polarized part of the network. Based on this intuition, we continue with our method which is separated into two main components: the offline part and the online part.
Offline
This part of our method refers to the process of training the GCN to predict the decrease of the polarization after setting a node’s external opinion to zero. Let:
$Gain(v)=\pi(\mathbf{z^{ss}_{old}})-\pi(\mathbf{z^{ss}_{new}})$
where $\mathbf{z^{ss}_{old}}$ and $\mathbf{z^{ss}_{new}}$ are the steady state vectors of external opinions before and after setting $z_{v}=0$ respectively.
In other words, we want to train a GCN to predict the $Gain$ for each node. This is a node regression task because it predicts a continuous target value. To be more precise, the GCN is responsible to generate the node embeddings. Then, those node embeddings are given as input to a regression head which outputs the predicted $Gain$. The regression head is usually a linear layer because the training is done in an end to end fashion and non-linearities are already applied in earlier stages of the pipeline (inside the GCN). We provide a detailed description of the data that were used for training in Section 6.
Online
Figure 1: Graphical illustration of the pipeline that finds the best node at a timestep
The online part of our method is what we actually use to solve the ModerateExpressed problem. It builds upon and aims to improve, in terms of time, the previously discussed GreedyExt algorithm. The core principle behind both algorithms is iterative: at each timestep, we aim to select the best node to add to the solution set based on its contribution to the reduction of polarization in the network. However, our approach introduces a key innovation that dramatically enhances efficiency and scalability.
The fundamental difference lies in how the “best” node is identified. In the original GreedyExt algorithm, this process involved a brute-force search where, at each step, the polarization of the network was computed after setting the opinion of each node to zero individually. This required recalculating polarization multiple times, leading to substantial computational overhead, especially for large graphs. In contrast, our algorithm replaces this brute-force search with a trained neural network model, which predicts the polarization reduction directly for each node. Figure 1 summarizes the pipeline for finding the best node.
Breifly, the Graph is given as input to the GCN, which constructs one embedding for each node. Then the embeddings are fed to a regression head (a linear layer), which outputs one value per embedding. Finally, with argmax we locate the node which is the best at this timestep, according to our model.
In summary, by integrating machine learning into the node selection process, our algorithm aims to significantly improve upon the GreedyExt method, in terms of time required to construct a solution set, that is almost as good as the one that GreedyExt finds. It maintains the iterative structure, but replaces brute-force computation with a more scalable approach, reducing complexity, while preserving accuracy in the identification of key nodes. We refer to this algorithm as the GNN-GreedyExt in the experiments that follow. Algorithm 2 contains the pseudocode for the GNN-GreedyExt.
Algorithm 2 GNN-GreedyExt
  Input: Graph $G=(V,E)$, $\mathbf{s}$, $\mathbf{z}$, Maximum timesteps $K$, Pre-trained GNN model
  Initialize solution set $T\leftarrow\emptyset$
  for $t=1$ to $K$ do
[…]
     Select node $v^{*}\leftarrow\arg\max_{v}\textit{scores}(v)$
      $T\leftarrow T\ \cup\ \left\{v^{*}\right\}$
     Update $G\leftarrow$ set opinion of $v^{*}$ to 00 in $G$
     Compute the new equilibrium and isolate node $v^{*}$
  end for
  Output: Solution set $T$
6 Datasets and Training
We used both synthetic and real networks, which we present in this section. Realistic synthetic networks are required in order to facilitate the training process of the GCN [^4] It was in general not easy to find hydrated instances of social networks with two echo chambers and ground truth communities (red blue). Due to this limited availability of data we decided to train on artificial networks which try to capture some important properties of real networks, while real networks are necessary to evaluate our method on real scenarios.
Degree Corrected Stochastic Block Model
Earlier we discussed about the node regression task, which we commonly refer to as the gain regression task. In order to train the GCN for this task we need graphs where each node is accompanied with a target value that our model has to learn to predict. This value is no other than the gain that results from moderating the node (setting it’s expressed opinion to zero). As a first step towards this direction, we construct synthetic graphs with two echo chambers. The graphs are generated using a degree corrected stochastic block model (DCSBM), which was introduced in (Karrer & Newman, 2011). The difference of this model with other stochastic block models lies in the degree distribution. In the DCSBM model the degree distribution of the graph follows a power law. The details of the DCSBM model can be found in the paper (Karrer & Newman, 2011).
Figure 2: Example of graph generated with the DCSBM model.Node size reflects degree, with larger nodes having higher degrees. Blue and red nodes represent two echo chambers.
We construct 128 graphs with the DCSBM model, where each graph consists of 1000 nodes. The degrees are drawn from a power law distribution, which is common in social networks according to (Newman, 2010). Next we assign internal opinions to nodes based on their community membership. Nodes $v$ in community 1 are set $s_{v}=+1$ while nodes $u$ in community 2 are set $s_{u}=-1$. Finally, $\mathbf{z^{ss}}$ is computed and $z_{i}^{ss}$ is attached to each node $i$. As a result every node consists of two features: $s_{i}$, and $z_{i}$.
As we mentioned earlier, training for the node regression task requires the actual gain resulting from the moderation of each node, as supervision. This is computed as a preprocessing step of our graph data.
Regarding the precise GCN architecture, we used a 2-layer GCN with 16 hidden units per layer resulting in 16 dimensional node embeddings. A linear layer with 16 input values and one output value follows to transform the node embeddings into predicted gains.
This is a relatively simple model with a small number of parameters, but it works very well in practice. Although more complex models with a large number of parameters perform slightly better in the node regression task, when it comes to inference (solve ModerateExpressed problem) we observe no difference in the solutions. This is probably because the inference (online) task depends on the ordering of the predicted nodes’ gains rather than their exact values. Simply put, despite the predicted gains of the simple model (16 dim - 2 layer GCN) being less accurate compared to complex models, the ordering of the simple model’s predictions is consistent with the ordering of the predictions of complex models (i.e. if node A is better than B based on the complex models’ predictions, the same holds for the simple model). For this reason the inferred nodes are the same in both cases.
Finally, a simpler model has an extra advantage in terms of the time required for the algorithm to complete compared to scenarios with more complex models. This is because a smaller neural network results in faster forward passes. As a result, for our experiments we used this simple GCN with 2 layers where each layer consisted of 16 hidden units.
Real Networks
We evaluated the GNN-GreedyExt in both real and synthetic networks. Here we present the real networks that we used. The experimental results are discussed in the next section. Some important properties characterizing both synthetic and real networks are summarized in Table 1
Graph $|V|$ $|E|$ $\pi$
Books 105 441 0.107
LiveJournal 2766 24138 0.179
WireTaping 5000 201876 0.011
DCSBM 977 3724 0.102
Table 1: Characteristics of the real networks considered. The last row refers to a synthetic (DCSBM-generated) graph that was used for training
Political Books
This dataset consists of political books that were sold on Amazon. The edges between political books represent frequent co-purchasing. The books are labeled as Conservative, Liberal and Neutral. The opinion mapping that we use is the following: Conservative: 1, Neutral: 0, Liberal: -1. The reason why we choose these 3 values is because we have no means to quantify numerically for each node what their exact opinion is. One alternative is to randomly sample opinions in the negative side of the spectrum of Liberal nodes, and in the positive side of the spectrum for the Conservative nodes. Finally, for the neutral nodes we could sample opinions around zero. There is no big difference in terms of our methodology. Of course, the resulting polarisation of the network and the final selected nodes by the algorithms would be different as their selection or not from the algorithms, depends both on their opinion and graph structure. But we are examining whether the GNN-GreedyExt that we propose makes equivalent choices to the GreedyExt algorithm, so the starting point of the network doesn’t really make any difference. In both cases the GreedyExt algorithm will make its greedy choices, and our algorithm should present similar behavior. Thus, without loss of generality, we stick with the +1,0, -1 mapping that we mentioned which is consistent with (Matakos et al., 2017) [^5] This opinion mapping was used for this dataset in their work and facilitates comparison with their results. The polarisation for the Political Books dataset is $\pi=0.107$. It consists of 105 nodes and 441 edges.
LiveJournal
LiveJournal is a free online blogging community where users declare friendship to each other. LiveJournal also allows users to form a group, to which other members can then join. Such user-defined groups are considered as ground-truth communities. This dataset was introduced in (Yang & Leskovec, 2012) and is made public from the authors (SNAP Datasets). It consist of multiple communities. Because of our interest in the dual echo chamber structure, a network with two communities was required. For this reason, we started with the 5000 best communities according to the original paper (Yang & Leskovec, 2012), where they define what makes a community “good”. Then we sorted them based on their length and constructed the graph for the two largest communities. The resulting graph is depicted in Figure 3 and clearly has the dual echo chamber structure. The opinions are assigned based on community membership.
Figure 3: LiveJournal Graph
Twitter Wiretaping Scandal
Figure 4: Retweet graph for the wiretaping scandal with 5k nodes
This dataset contains a snapshot of the political discussion on Twitter regarding the Greek Wiretaping Scandal, which was revealed in 2022. The detailed description of this dataset is given in (Dimitriadis et al., 2024). Briefly, it contains multiple tweets with hashtags and keywords relevant to the scandal. Using this dataset we construct the retweet graph. To infer user opinions, we adopt a straightforward heuristic. We count the number of politicians that a user follows and assume that the party she supports is the one that she follows the most politicians from. In case of a draw we assume she doesn’t support any party. Finally, we set her opinion to $s_{i}=+1$ if she supports a party that is left on the political spectrum and $s_{i}=-1$ otherwise. Because the retweet graph consists of 44k nodes, we work with a smaller sample of 5k nodes which is depicted in Figure 4. Due to its high density, this graph doesn’t present as high polarization as the other graphs.
7 Experiments
In this section we present the experimental results of our algorithm and compare it to the GreedyExt. For all experiments we choose small values of $K$ (10% of the graph’s size)
In order to ensure that the problem instance we are dealing with is not “trivial”, we also implement a random baseline algorithm, that picks the $K$ nodes to moderate uniformly. Clearly, scenarios in which the random scheme is competitive to the GreedyExt one suggest that either (i) this scenario has not enough social “structure” with all nodes being more or less equivalent (in their influence role) and/or (ii) $K$ is too large for this scenario [^6] It is easy to see that as K increases, the problem of “correctly” picking nodes becomes easier.
Figure 5: Polarization vs K for the Books Dataset
De-polariszation achieved
In this first set of experiments, we use one of our real datasets (the Politcal Books one) to demonstrate these two key points: (i) the discussed impact of $K$, but that the problem is not “trivial” for a range of such values, even in a small scenario, and (ii) that our scheme’s performance is close to the “optimal” GreedyExt one.
In Figure 5 we show the final polarization index of Equation 3, on the y axis, as the number of nodes $K$ to moderate increases, for three algorithms: (i) GreedyExt, (ii) GNN-GreedyExt and (iii) the random baseline.
Key Observations
  1. 1.
    The GNN-GreedyExt algorithm follows GreedyExt closely, which means that it makes equivalent choices at each timestep.
  2. 2.
    The random baseline algorithm fails to decrease the polarization as much as the other algorithms, a fact which implies that the problem is not trivial
Similarly, we evaluate our method on the other real networks. The results of our experiments for the real networks are summarized in Figure 6. According to this figure, the GNN-GreedyExt always achieves similar polarization to GreedyExt. The orange bar is the initial polarization for each dataset, while the blue and the green bars show the polarization achieved from GreedyExt and GNN-GreedyExt respectively, after setting $K$ nodes to zero. Precisely, the polarization decrease of GNN-GreedyExt is roughly 95% of that achieved by GreedyExt.
Figure 6: Polarization of each algorithm per dataset. The orange bars show the initial polarization of each network. The blue and green bars show the final polarization that GreedyExt and GNN-GreedyExt achieve
Inference speed
So far, we have only discussed how the two algorithms compare, in terms of the final polarization that they achieve. As we have already mentioned, the GreedyExt algorithm, is not efficient for large graphs, because of its high computational complexity. Particularly, the evaluation of every node independently at each timestep increases the complexity dramatically. This is exactly where we expect the GNN-GreedyExt to excel, achieving the demonstrated close to optimal depolarization much faster. In the following set of experiments we demonstrate two key points: (i) the GNN-GreedyExt outperforms the GreedyExt in terms of inference speed, and (ii) the inference speed advantage of GNN-GreedyExt increases as the depolarizing graph’s size increases.
In Figure 7 the two algorithms are compared over the LiveJournal and Wiretaping datasets. Each bar corresponds to the time required (in seconds) for the corresponding algorithm to complete. For this experiment we set $K=50$ without loss of generality.
Figure 7: Required time per algorithm for the LiveJournal and Wiretaping datasets
As the picture shows, the GNN-GreedyExt outperforms the GreedyExt in terms of time. Similar results hold for the other datasets as well.
Finally, the next experiment compares the time to complete of the two algorithms for graphs with various sizes. In Figure 8 the y-axis shows the required time to complete, while the x-axis has the graphs’ sizes. As we can see, the required time for the GreedyExt algorithm grows very quickly with respect to the graph size. In contrast the time to complete for the GNN-GreedyExt grows much slower, with respect to the size of the graph.
Figure 8: Required time per algorithm for graphs with various sizes
Moreover, the time advantage of the GNN-GreedyExt against the GreedyExt, is larger, for larger graphs. In other words, it increases with respect to the size of the underlying graph. In the examples that we have shown where large graphs were involved, the LiveJournal and the Wiretaping datasets, the GNN-GreedyExt is ×$\times 10$ and ×$\times 16$ times faster respectively.
8 Conclusion
In this work, we proposed an alternative approach to address the ModerateExpressed problem, originally introduced in (Matakos et al., 2017). One of the key limitations of their solution is its lack of scalability, as their algorithm becomes increasingly slow, when applied to larger graphs. Our method leverages Graph Neural Networks (GNNs), specifically a 2-layer Graph Convolutional Network (GCN), to generate node embeddings. Instead of using brute force evaluation, we apply a simple argmax function to identify the node with the highest score. In terms of polarization, our approach achieves results almost identical to those of the original algorithm. However, the main advantage of our method is its ability to scale efficiently to larger graphs due to the GNN architecture. Our results show that the GNN-GreedyExt dramatically improves the required time to solve the problem and thus it is a great choice when dealing with large graphs. An interesting future direction is to explore the submodularity of the problem ModerateExpressed and quantify the suboptimality for greedy algorithms like the ones we have discussed.
Acknowledgements
The authors have been supported in part by the Hellenic Foundation for Research & Innovation (HFRI) project 8017: “AI4RecNets: Artificial Intelligence (AI) Driven Co-design of Recommendation and Networking Algorithms”